OpenStudy (anonymous):

At the bookstore, 40 copies of a popular paperback book fit perfectly on a 5-foot shelf. How many copies of this book would fit perfectly on an 8-foot shelf?

OpenStudy (tkhunny):

Looks like a job for a proportion. Be careful not to count fractional books. The wording of the problem suggests you won't encounter any fractional books, but I thought I'd mention it for future reference.

OpenStudy (anonymous):

Divide 40 by 5. That equals 8 books per ft. Then multiply 8 by 8.

OpenStudy (anonymous):

64?

OpenStudy (tkhunny):

Why do you doubt? Did it come out exact or do you have to cut up a book?

OpenStudy (anonymous):

damn no need to be rude...

OpenStudy (tkhunny):

1) I'm never rude. 2) "Why do you doubt." is encouragement for you to develop independent thinking. 3) "...cut up a book?" is just a question. Feel free to answer it.

OpenStudy (anonymous):

dang hunny it was a joke ... havent u heard the song ...

OpenStudy (tkhunny):

That's fine. Are you going to answer the question?

OpenStudy (anonymous):

ill answer... no i ddnt have to cut up a book lol and i only doubt because i did it in my head

OpenStudy (tkhunny):

I'm not talking about the calculation. Did the answer come out EXACTLY 64, or was it 63.99999842 = just a hair short of 64?

OpenStudy (anonymous):

ohh no i acutally got 62 but 64 was the closest answer

OpenStudy (tkhunny):

You didn't listen to kasie. \(\dfrac{40}{5} = \dfrac{What}{8}\)??

OpenStudy (anonymous):

i did 40/5 n got 8 then i did 8x8=64

OpenStudy (tkhunny):

Okay, that's enough of this one. Move on to the next one.

OpenStudy (anonymous):

k